OpenStudy (anonymous):

What is 3 to the power of 2 over 3 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

\[3^{\frac{ 2 }{ 3 }} =(3^2)^{\frac{ 1 }{ 3 }}=9^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

which answer choice

imqwerty (imqwerty):

A :D